# LLM-as-a-Judge: Reassessing the Performance of LLMs in Extractive QA

### v1

@@ Abstract
Extractive reading comprehension question answering (QA) datasets are typically evaluated using Exact Match (EM) and F1-score, but these metrics often fail to fully capture model performance. With the success of large language models (LLMs), they have been employed in various tasks, including serving as judges (LLM-as-a-judge). In this paper, we reassess the performance of QA models using LLM-as-a-judge across four reading comprehension QA datasets. We examine different families of LLMs and various answer types to evaluate the effectiveness of LLM-as-a-judge in these tasks. Our results show that LLM-as-a-judge is highly correlated with human judgments and can replace traditional EM/F1 metrics. By using LLM-as-a-judge, the correlation with human judgments improves significantly, from 0.17 (EM) and 0.36 (F1-score) to 0.85. These findings confirm that EM and F1 metrics underestimate the true performance of the QA models. While LLM-as-a-judge is not perfect for more difficult answer types (e.g., job), it still outperforms EM/F1, and we observe no bias issues, such as self-preference, when the same model is used for both the QA and judgment tasks. [^1] Our data and code are available at https://github.com/Alab-NII/llm-judge-extract-qa
@@ 1 Introduction
Our results show that using LLM-as-a-judge is highly correlated with human judgments, improving from 0.17 (EM) and 0.36 (F1-score) to 0.85, highlighting its potential to replace EM and F1 metrics. This high correlation further confirms that EM and F1 metrics underestimate the true performance of the models. Our analysis demonstrates that LLMs as judges are particularly effective for answer types related to numbers and dates in extractive QA. Moreover, we observe no self-preference bias in the extractive QA task when the same model is used for both the QA and judgment tasks.
@@ 4.1 Reliability of LLM-as-a-Judge Scores
Qwen 2 7B 0.598 0.781 0.863 0.198 0.390
Qwen 2 72B 0.653 0.680 0.793 0.145 0.259
Gemma 2 9B 0.715 0.833 0.848 0.242 0.488
Gemma 2 27B 0.711 0.850 0.908 0.210 0.408
Llama 3.1 8B 0.700 0.876 0.945 0.271 0.521
Llama 3.1 70B 0.626 0.803 0.848 0.040 0.257
Average 0.653 0.750 0.847 0.174 0.359
Table 2: Pearson correlation coefficients between human judgments and the LLM-as-a-judge evaluations from three models—Mistral 7B v0.3, Llama 3.3 70B, and Qwen 2.5 72B—across eight different QA models are presented. Additionally, we include the correlation scores between human judgments and the EM/F1 scores.

### v2

@@ Abstract
Extractive reading comprehension question answering (QA) datasets are typically evaluated using Exact Match (EM) and F1-score, but these metrics often fail to fully capture model performance. With the success of large language models (LLMs), they have been employed in various tasks, including serving as judges (LLM-as-a-judge). In this paper, we reassess the performance of QA models using LLM-as-a-judge across four reading comprehension QA datasets. We examine different families of LLMs and various answer types to evaluate the effectiveness of LLM-as-a-judge in these tasks. Our results show that LLM-as-a-judge is highly correlated with human judgments and can replace traditional EM/F1 metrics. By using LLM-as-a-judge, the correlation with human judgments improves significantly, from 0.22 (EM) and 0.40 (F1-score) to 0.85. These findings confirm that EM and F1 metrics underestimate the true performance of the QA models. While LLM-as-a-judge is not perfect for more difficult answer types (e.g., job), it still outperforms EM/F1, and we observe no bias issues, such as self-preference, when the same model is used for both the QA and judgment tasks. [^1] Our data and code are available at https://github.com/Alab-NII/llm-judge-extract-qa
@@ 1 Introduction
Our results show that using LLM-as-a-judge is highly correlated with human judgments, improving from 0.22 (EM) and 0.40 (F1-score) to 0.85, highlighting its potential to replace EM and F1 metrics. This high correlation further confirms that EM and F1 metrics underestimate the true performance of the models. Our analysis demonstrates that LLMs as judges are particularly effective for answer types related to numbers and dates in extractive QA. Moreover, we observe no self-preference bias in the extractive QA task when the same model is used for both the QA and judgment tasks.
@@ 4.1 Reliability of LLM-as-a-Judge Scores
Qwen 2 7B 0.598 0.781 0.863 0.234 0.434
Qwen 2 72B 0.653 0.680 0.793 0.157 0.274
Gemma 2 9B 0.715 0.833 0.848 0.288 0.572
Gemma 2 27B 0.711 0.850 0.908 0.247 0.487
Llama 3.1 8B 0.700 0.876 0.945 0.331 0.628
Llama 3.1 70B 0.626 0.803 0.848 NaN 0.281
Average 0.653 0.750 0.847 0.220 0.404
Table 2: Pearson correlation coefficients between human judgments and the LLM-as-a-judge evaluations from three models—Mistral 7B v0.3, Llama 3.3 70B, and Qwen 2.5 72B—across eight different QA models are presented. Additionally, we include the correlation scores between human judgments and the EM/F1 scores. NaN values arise when none of the predicted answers by Llama 3.1 70B exactly match the gold answers, resulting in a constant list of 0s, for which correlation is undefined. This often occurs when gold answers are plain numbers (e.g., 93.5), but predictions include extra context like percentages (e.g., 93.5%) or time spans (e.g., 38 years).